Sandy’s Sweatshirt Shop sells college sweatshirts. White sweatshirts sell for $18.95 each and red ones sell for $19.50 each. If receipts for the sale of 30 sweatshirts total $572.90, how many of each color did the shop sell? Complete the following table, letting the number of white sweatshirts and the number of red sweatshirts.
Answer by richwmiller(17219)

Of course, since there is no table included we can't complete it.
let w=white sweatshirts and r=red sweatshirts
30 sweatshirts were sold
r+s=30
s1895+r1950=57290
r=30-s
substitute 30-s for r
s*1895+(30-s)*1950=57290
s*1895+30*1950-s1950=57290
s*1895-s1950=57290-30*1950
Dig out the calculator.
30*1950=
57290-30*1950=
s1895-s1950=
Don't worry about the minus signs on both sides.
they go away by dividing
minus divided by minus= +